# Wake Vision: A Tailored Dataset and Benchmark Suite for TinyML Computer Vision Applications

### v3

@@ Appendix H Dataset Access and Organization
Wake Vision is available online via TFDS and Hugging Face Datasets. We exclude direct links to the dataset during the review period to avoid breaking anonymity.
@@ Appendix I Label Generation Details
To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions. Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Sec. 3.3.
@@
The full Open Images v7 dataset is not hosted by the dataset authors. Rather it is provided as a collection of flickr image Uniform Resource Locators (URLs) and their associated labels. As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform. As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Tab. 7.
@@
The following labels are considered body parts and are labelled according to the body part configuration.:
@@
To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections. These label corrections are described in Sec. 3.2 and deprecate the automatic label corrections described in this section.
@@
The DOI and URLs will be provided after the review period to avoid breaking anonymity.
Appendix S Datasheet for Wake Vision
@@
These details will be provided after the review period to avoid breaking anonymity.
@@
These details will be provided after the review period to avoid breaking anonymity.
@@
Yes, we open-source our filtering code. https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
@@
Yes, the estimated error rates are reported in Table 2. We also discuss the high likelihood of false negatives in Wake Vision’s training sets in Section 3.2. These errors are due to the labeling and filtering process. We mitigate the impact of these errors by manually labeling the validation and test sets as well as study their impact for TinyML models in Section 4.4.
@@
Removed during review to maintain anonymity.
@@
Yes, we will continue to host older versions of the dataset. Details will be provided after review to preserve anonymity.

### v5

@@ Appendix H Dataset Access and Organization
Wake Vision is available online at the Harvard Dataverse, via TFDS and Hugging Face Datasets. The version on Harvard Dataverse has been assigned the following DOI: https://doi.org/10.7910/DVN/1HOPXC More information can be found on the https://wakevision.ai webpage.
@@ Appendix I Label Generation Details
To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions. Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Section 3.3.
@@
The full Open Images v7 dataset is not hosted by the dataset authors. Rather it is provided as a collection of flickr image Uniform Resource Locators (URLs) and their associated labels. As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform. As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Table 7.
@@
The following labels are considered body parts and are labelled according to the body part configuration:
@@
To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections. These label corrections are described in Section 3.2 and deprecate the automatic label corrections described in this section.
@@
The version of Wake Vision on Harvard Dataverse has been assigned the following DOI: https://doi.org/10.7910/DVN/1HOPXC The following two croissant files describe the Wake Vision (Quality, Validation, and Test) and Wake Vision (Large) respectively: https://huggingface.co/api/datasets/Harvard-Edge/Wake-Vision/croissant https://huggingface.co/api/datasets/Harvard-Edge/Wake-Vision-Train-Large/croissant
Appendix S Car and Bird Datasets
Table 10 compares the number of images as well as the label error rate for the car and bird datasets, when generated using both our methodology and the VWW one. All datasets are balanced, i.e., have the same number of images for the target and background class. Label error rates have been manually computed on a subset of 500 randomly selected images for each generated dataset. Examples of positive class images from the Wake Vision test splits are shown in Figure 11
Table 10: Raw data of Car and Bird datasets.
Methodology # of Images Label Error Rate
Car
Wake Vision (Ours) 157,668 4.8%
VWW [7] 12,902 6.4%
Bird
Wake Vision (Ours) 31,642 0.6%
VWW [7] 1,148 6.6%
(a) Wake Vision “Car” samples
(b) Wake Vision “Bird” samples
Figure 11: Positive class examples from the test splits of our Wake Vision generated datasets for “Car” and “Bird” objects, demonstrating our ability to easily prototype large-scale TinyML vision datasets (Section 3.5).
Appendix T Wake Vision Challenge Results
Submissions to the model-centric track have been trained using the same recipe adopted in Section 4.3. Table 11 provides input shape, Flash, RAM, and MACs for each model presented in Section 4.5 as well as the mean and standard deviation of test accuracies for the Wake Vision dataset. RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”.
Table 11: Raw data of model-centric track.
Model input RAM Flash MACs Test Acc
.tflite volume [kiB] [kiB] [MM] [%]
anas-benalla (50,50,3) 48.5 104.55 4,899,292 $79.7\pm 0.28$
samy (80,80,3) 73.5 34.55 5,718,064 $79.5\pm 0.61$
benx13 (48,48,3) 63.5 94.95 1,693,474 $79.6\pm 0.22$
cezar (50,50,3) 245.5 57.23 20,435,868 $79.6\pm 0.42$
mohammad_hallaq (80,80,3) 129.5 128.32 3,887,331 $77.3\pm 0.5$
ymac (50,50,3) 34.5 27.77 431,985 $76.7\pm 0.51$
apighetti (50,50,3) 48.5 278.36 693,818 $74.3\pm 74.3$
Table 12 shows test accuracy, precision, and recall of a MobileNetV2 having a width modifier of 0.25, trained on the original Wake Vision large split and on the same split enhanced with the technique proposed by the winner of the data-centric track [4]. The training recipe is the same reported in Section 4.3.
Table 12: Main improvement of the data-centric track.
Metrics Train Large Split
Original Enhanced
MobileNetV2_0.25
Test Accuracy $82.3\pm 0.2$ $87.1\pm 0.81$
Precision $86.4\pm 1.49$ $87.6\pm 1.78$
Recall $76.7\pm 1.83$ $86.5\pm 1.32$
Appendix U Training Resources
A desktop computer featuring a 13th Gen Intel® Core™ i9-13900K × 32, a 32 GB of RAM, and an NVIDIA GeForce RTX 4080 with 16 GB of VRAM has been employed to train the TinyML models.
Appendix V Datasheet for Wake Vision
@@
The authors of this paper created the dataset (Harvard Edge Computing Lab, Technical University of Denmark, and Useful Sensors).
@@
The Google TPU Research Cloud program partially supported this work via cloud computing credits. The PhD stipend of Emil Njor was supported by the Innovation Fund Denmark DIREC project (9142-00001B). Emil Njor’s external research stay was furthermore supported by Fulbright Denmark, Stibo Fonden, Thomas. B. Thriges Fond, Otto Mønsteds Fond and Kaj og Hermilla Ostenfelds Fond. This work was also support by the National Science Foundation (NSF) and the Semiconductor Research Corporation (SRC).
@@
Yes, we open-source our filtering code. https://github.com/harvard-edge/Wake_Vision
@@
Yes, the estimated error rates are reported in Table 2. We also discuss the high likelihood of false negatives in Wake Vision’s training sets in Section 3.2. These errors are due to the labeling and filtering process. We mitigate the impact of these errors by manually labeling the validation and test sets as well as study their impact for TinyML models in Section C.
@@
At the time of writing, the authors can be contacted at cbanbury@g.harvard.edu. Once maintenance is transferred to the Edge AI Foundation, there will be a point of contact there.
@@
Yes, we will continue to host older versions of the dataset. This is an inherent feature of the Harvard Dataverse.